OpenStudy (anonymous):

Simplify the expression: 10(y^5/x^2)^2

OpenStudy (anonymous):

\[\huge = 10 (\frac{y^{10}}{x^4})\]

OpenStudy (anonymous):

\[\huge (a^b)^c = a^{b \times c}\]

OpenStudy (anonymous):

power of a power you multiply,

OpenStudy (anonymous):

hehe, waterineys show it better gl

OpenStudy (anonymous):

thanks guys!!